OpenStudy (anonymous):

subtract polynomials solve please (7.2x-3.4y+z)(-9.2x+4.8y-2.1z) solve please

OpenStudy (anonymous):

First you distribute the understood 1 to both parenthesis so you can get rid of the parenthesis. 7.2x - 3.4y + z - 9.2x + 4.8y - 2.1z Combine Like Terms 7.2x - 9.2x = -2x -3.4y + 4.8y = 1.4y z - 2.1z = -1.1z Now put them together -2x + 1.4y - 1.1z